OpenStudy (anonymous):

i have this word problem. the second of two numbers is 6 times the first. their sum is 77. find the numbers

OpenStudy (anonymous):

Each of those sentences is an equation. Let h be the first number, and k be the second. Now write the two equations

OpenStudy (anonymous):

thats what i don't understande

OpenStudy (anonymous):

x+6x=77

OpenStudy (anonymous):

Yes, that's right.

OpenStudy (anonymous):

Or as I was going to suggest, k = 6h k+h = 77 So 6h + h = 77 We can add six h's to one h and get 7h = 77 Which means h =?